OpenStudy (anonymous):

The Nygen family wants to put a circular fence around their pool. They want it to be exactly 12 feet from the center of the pool. How many feet of fencing do they need to buy?

OpenStudy (anonymous):

\[\pi (2*12)=24 \pi =75.4 \text{ feet} \]